### v1

# MCPZoo: A Large-Scale Dataset of Runnable Model Context Protocol Servers for AI Agent

Abstract
Model Context Protocol (MCP) enables agents to interact with external tools, yet empirical research on MCP is hindered by the lack of large-scale, accessible datasets. We present MCPZoo, the largest and most comprehensive dataset of MCP servers collected from multiple public sources, comprising 90,146 servers. MCPZoo includes over ten thousand server instances that have been deployed and verified as runnable and interactable, supporting realistic experimentation beyond static analysis. The dataset provides unified metadata and access interfaces, enabling systematic exploration and interaction without manual deployment effort. MCPZoo is released as an open and accessible resource to support research on MCP-based security analysis.
Index Terms:
Model Context Protocol, MCP, Agents, Agent Security, Security Measurement
1 Introduction
The software ecosystem is witnessing a paradigm shift driven by Large Language Models (LLMs) and autonomous Agents. Just as mobile applications revolutionized digital interaction over the last decade, "Tools"—interfaces that allow Agents to perceive and manipulate the world—are becoming the fundamental building blocks of the AI era. The Model Context Protocol (MCP) has rapidly emerged as the de facto standard for connecting these AI models with diverse data sources and environments.
However, conducting large-scale empirical research on this growing ecosystem presents significant challenges. While recent studies have begun to explore the security risks and robustness of MCP integrations (e.g., prompt injection vulnerabilities in tool use), these efforts are often constrained by the difficulty of setting up testing environments. Unlike mobile apps which are packaged as standalone binaries, MCP servers are diverse software projects with complex, often conflicting dependencies. Consequently, existing research is frequently limited to small, manually curated datasets, leading to experiments that are difficult to reproduce or generalize.
Inspired by AndroZoo [androzoo], which democratized Android research by collecting millions of apps for the community, we present MCPZoo. To the best of our knowledge, MCPZoo is the largest collection of Model Context Protocol servers to date. While AndroZoo solved the problem of access, MCPZoo aims to solve the problem of runnability.
By providing a standardized, “unrestricted, and scalable” dataset of runnable agents, we aim to enable the research community to move beyond small-scale testing and conduct comprehensive, reproducible analyses on the security and capability of the AI agent ecosystem.
Contribution. MCPZoo makes the following contributions:
  • The largest and most comprehensive dataset of MCP servers currently available, with 90,146 MCP servers.
  • The first dataset to include more than ten thousand runnable and interactable server instances.
  • Released as an open and accessible dataset to support community-driven MCP analysis.
2 Build Infrastructure
2.1 Construct MCP Zoo
Sources. We collected MCP server projects from six public sources: MCP World [mcpworld], MCP Store [mcpstore], MCP Servers Repository [mcprepository], AIbase MCP [aibasemcp], Mcp.so [mcpso], and Pulse MCP [pulsemcp]. Each source provides an independently curated directory of MCP servers, reflecting different community and platform perspectives.
Data Fields. Both metadata and source code for each MCP server are saved. We extract a unified set of metadata fields, including the server name, description, overview text (detailed Readme), source code URL, author information, etc. In parallel, we download the corresponding code repositories and organize all artifacts in a structured storage layout. To support cross-source alignment and efficient access, we generate symbolic links that associate multiple listings with their code snapshots.
De-Duplication. To ensure high data quality, we de-duplicate large numbers of forks, mirrors, and copy-paste repositories. Our de-duplication process follows three steps. We first normalize repository URLs and clean metadata to resolve redirects and naming inconsistencies. We then vectorize representative textual content extracted from configuration and code files to capture semantic similarity between server implementations. Finally, we cluster highly similar servers and retain a single representative per cluster, while preserving provenance information for all sources.
2.2 Make the Zoo Alive
Building Docker Image. To make servers runnable at scale, we employ an automated build agent to convert the MCP server source code into executable Docker images. The agent performs context-aware analysis of code and dependencies to infer runtime settings and generate Docker configurations. When build failures occur, it iteratively analyzes error signals and regenerates the Docker files with targeted adjustments until the image builds successfully or a retry limit is reached.
Liveness Check. A successfully built container does not necessarily indicate a runnable MCP server. We therefore perform a standardized liveness check to validate whether a server is practically usable in the MCP setting. This check consists of two complementary aspects. The first aspect examines execution capability, accounting for common failure cases such as missing configurations or hardcoded paths. The second aspect examines protocol-level functionality through basic interactions, including retrieving declared tool lists. Only servers that pass both execution and interaction checks are considered alive in MCPZoo.
3 MCPZoo
3.1 Overview
MCPZoo is a large-scale, continuously growing dataset of Model Context Protocol (MCP) servers collected from multiple public sources. At the time of writing, MCPZoo contains 90,146 distinct MCP servers, whose source code collectively amounts to 377 GB, among which 14,206 servers are verified to be runnable. Table I summarizes the composition of MCPZoo by source, reporting the number of collected servers from each of the platforms.
TABLE I: Statistics of MCP Servers from Six Public Sources
Source # Collected Percentage
MMCP World 31,048 34.44%
MCP Store 25,695 28.50%
MCP Repository 14,341 15.91%
AIbase MCP 11,120 12.33%
Mcp.so 6,772 7.51%
Pulse MCP 1,170 1.30%
Total 90,146 100.00%
The dataset is constructed through an ongoing collection effort and is actively expanding as new MCP servers emerge in the ecosystem.
3.2 Visit MCPZoo
MCPZoo is publicly accessible through a website [^1] https://security.fudan.edu.cn/zoo. The website provides downloadable metadata covering the entire collection of MCP servers, enabling researchers to systematically inspect and filter the dataset at scale. A summary of all available metadata fields is presented in Table II.
TABLE II: Metadata fields provided by MCPZoo
Field Description
serverName The human-readable name of the MCP server as provided by the source platform.
description A short textual description summarizing the functionality and purpose of the server.
source The originating platform or repository from which the MCP server was collected.
creator The declared creator or organization of the MCP server.
overview A more detailed overview of the server, typically describing its supported tools, usage scenarios, or design intent.
MCPZoo provides the capability to dynamically invoke MCP servers through remote connections. At present, 14,206 servers have been deployed, verified, and confirmed to be interactable in real execution environments. For each server, MCPZoo maintains a unified remote access configuration, which allows users to directly connect to and interact with MCP servers in a consistent manner. This design lowers the barrier to testing and experimentation. The MCPZoo website publicly exposes access interfaces for 15 runnable MCP servers for trial use. Researchers who require access to a larger set of interactable servers may contact us for extended access.
3.3 Access Conditions
MCPZoo is made available to the research community to support measurement, systems, and security studies on the MCP ecosystem. We require that researchers requesting access to MCPZoo agree to the following conditions. (1) Legal Compliance. Researchers must assess the legal implications of downloading, storing, and analyzing the collected materials in accordance with applicable local laws and host institution policies. (2) No Redistribution. The dataset should not, in general, be redistributed or republished, either in full or in part. (3) Non-Commercial Use. The dataset must not be used for commercial purposes. (4) Responsible Interaction. Researchers are expected to interact with MCP servers in a responsible manner and to avoid disruptive, abusive, or harmful behaviors toward deployed services. Security testing beyond benign interaction or measurement should follow responsible disclosure principles and may require additional approval. (5) Faculty Endorsement. Access requests must be endorsed by a faculty member or an individual in a permanent research position, who agrees to these conditions and commits to responsible use of the dataset.
We kindly request that the use of MCPZoo be acknowledged, and encourage researchers to disclose the list of the MCP servers involved in their studies whenever feasible, to support reproducibility.
4 Leverageing MCPZoo
MCPZoo can be used as a shared experimental foundation for research on real-world MCP deployments. For example, it enables agent benchmarking by providing diverse, realistic tool environments for evaluating agents’ planning and tool-use capabilities. It can also support security analysis at scale, such as systematically examining MCP servers for prompt injection or sandbox escape risks, extending prior studies on LLM and agent security to a protocol-level ecosystem. In addition, MCPZoo facilitates protocol evolution studies by allowing empirical analysis of MCP adoption and implementation patterns across a large set of real projects, in the spirit of prior large-scale software ecosystem measurements [androzoo].
5 Conclusion
We present MCPZoo, a large-scale dataset of MCP servers collected from multiple public sources, with a large number of servers verified as runnable and interactable. The dataset is publicly available, together with access interfaces, to support future research on MCP-based systems and analysis. MCPZoo can be accessed through the website: https://security.fudan.edu.cn/zoo.